OpenStudy (anonymous):

A horse is 45 pounds more than six times its rider's weight. If they weigh a total of 710 pounds, how much does the horse wiegh?

OpenStudy (anonymous):

6x + 45 + x = 710 7x = 710 - 45 7x = 665 x = 95 this is riders weight 6x + 45 570 + 45 615 this is the weith of the horse

OpenStudy (anonymous):

thanks soo much, :DDD <3

OpenStudy (anonymous):

no problem ^^